a packet of weight W is dropped with the help of a parachute and on striking the ground comes to reat with a retardation equal to twice the acceleration due to gravity. what is the force exerted on the ground?
  1. W
  2. 2W
  3. 3W
  4. 4W

Grade:11

1 Answers

VIGNESH VENKATESH
4 years ago
F=dp/dt
F=m(dv/dt)
Since W=mg, therefore m=W/g
dv/dt=2g (since acceleration)
THEREFORE, F=(W/g)*2g=2W